what is the three digit number that satisfies the following condition.the tens digit is greater than the ones digit, the sum of the digits is 9 and if digits are reversed and resulting three-digit number is subtracted from the original number, the difference is 198
Answer by Edwin McCravy(20081)

what is the three digit number that satisfies the following condition.the tens
digit is greater than the ones digit, the sum of the digits is 9 and if digits
are reversed and resulting three-digit number is subtracted from the original
number, the difference is 198
Let the number be ABC and its reverse be CBA,
Then the subtraction is:

 ABC
-CBA
 ---
 198

A must be bigger than C so that the top number will be bigger than the bottom
number.  So we will need to borrow 1 from the B at the start of the subtraction.

That will reduce the B in the top middle digit to 1 less than the B just below
it.

Therefore we will also need to borrow 1 from the A as well.

If we had not needed to borrow 1 from the A, then A-C would have been 1 to
get the 1 in 198.  But since we must borrow 1 from A, A-C must be 2
instead, so that borrowing 1 from A would give the 1 in 198.

A-C cannot be as high as 6-4 because all three digits must have sum 9,
and 6 and 4 alone have sum 10.  So we try lower digits.

We try A-C = 5-3, making the number 513 since the digits must have sum 9.
But that won't do because the tens digit is less than the ones digit.

We try A-C = 4-2. Then the number would be 432 since the digits must have sum 9.
That's one solution because the tens digit is greater than the ones digit, and
432-234=198.

We try A-C = 3-1. Then the number would be 351 since the digits must have sum 9. 
That's another solution because the tens digit is greater than the ones digit,
and 351-153 = 198.

So there are TWO solutions, 432 and 351

 432          351
-234         -153
----         ----
 198          198

Edwin